### v1

@@ 5 Proof of the main results from Section 3
Combining the above two bounds and since $h_{{\mathbf{a}}}\lesssim h_{F}$ for all $F\in\mathcal{F}_{{\mathbf{a}}}$ and $p\geq 1$, we infer that the assertion holds true.
@@ 6.1 Model problem
where ${\boldsymbol{J}}\in{{{\boldsymbol{H}}(\text{\rm div};D)}}$ is the source term. Both material properties $\boldsymbol{\epsilon}$} and $\boldsymbol{\nu}$} can vary in $D$ and are assumed to take positive-definite, symmetric, piecewise constant matrix values with eigenvalues uniformly bounded from above and below away from zero. Notice that, for simplicity, we are enforcing homogeneous Dirichlet boundary conditions (also called perfect electric conductor boundary conditions). The parameter $\omega$ scales as a frequency (reciprocal of a time scale) and can be either a real number (leading to the so-called definite Maxwell problem) or a pure imaginary umber (leading to the so-called indefinite problem). Whenever $\omega=0$, the additional property $\nabla{\cdot}{\boldsymbol{J}}=0$ is prescribed. For the time being, we assume that $\omega$ is a nonzero real number and refer the reader to Remarks 6.2 and 6.10 for further insight into the other cases.
@@ 6.2 Symmetric interior penalty dG discretization
For all $K\in\mathcal{T}_{h}$, we also define $\alpha^{\sharp}_{\partial K}$ so that $\alpha^{\sharp}_{\partial K}|_{F}:=\alpha^{\sharp}_{F}$ for all $F\in{\mathcal{F}}_{K}$. To avoid distracting technicalities, we assume that the material properties are only mildly heterogeneous and anisotropic. Thus, we hide anisotropy ratios such as $\alpha^{\sharp}_{K}/\alpha^{\flat}_{K}$ in the generic constants used in the error analysis. We do the same for contrast factors such as $\alpha^{\sharp}_{K}/\alpha^{\sharp}_{K^{\prime}}$, where $K^{\prime}$ lies in some neighborhood of $K$. We refer the reader to Remarks 6.1 and 6.4 for some further insight.
@@ 6.4 $hp$-a posteriori error analysis
Invoke the triangle inequality together with Lemmas 6.6 and 6.7, and recall the definition of $\eta$ from (6.13). ∎
@@
Local bounds on all terms composing the error estimator $\eta_{K}$ can be found from [12, Theorem 6.5] except for the term involving ${\nabla_{h}{\times}{\boldsymbol{E}}_{h}}]\!]_{F}^{\mathrm{d}}$. Invoking the $H^{1}$ to $L^{2}$ norm inverse inequality from [37, Theorem 4.76], we infer that

### v2

@@ 5 Proof of the main results from Section 3
Combining the above two bounds and since $h_{{\mathbf{a}}}\lesssim h_{F}$ for all $F\in\mathcal{F}_{{\mathbf{a}}}$ and $p\geq 1$, we infer that the assertion holds true. (Notice that we just dropped the factor $p^{-2\delta}$ scaling the tangential jumps of ${\boldsymbol{E}}_{h}$ since $\delta$ can be arbitrarily small.)
@@ 6.1 Model problem
where ${\boldsymbol{J}}\in{{{\boldsymbol{H}}(\text{\rm div};D)}}$ is the source term. Both material properties $\boldsymbol{\epsilon}$} and $\boldsymbol{\nu}$} can vary in $D$ and are assumed to take positive-definite, symmetric, piecewise constant matrix values with eigenvalues uniformly bounded from above and below away from zero. Notice that, for simplicity, we are enforcing homogeneous Dirichlet boundary conditions (also called perfect electric conductor boundary conditions). The parameter $\omega$ scales as a frequency (reciprocal of a time scale) and can be either a real number (leading to the so-called definite Maxwell problem) or a pure imaginary number (leading to the so-called indefinite problem). Whenever $\omega=0$, the additional property $\nabla{\cdot}{\boldsymbol{J}}=0$ is prescribed. For the time being, we assume that $\omega$ is a nonzero real number and refer the reader to Remarks 6.26 and 6.34 for further insight into the other cases.
@@ 6.2 Symmetric interior penalty dG discretization
For all $K\in\mathcal{T}_{h}$, we also define $\alpha^{\sharp}_{\partial K}$ so that $\alpha^{\sharp}_{\partial K}|_{F}:=\alpha^{\sharp}_{F}$ for all $F\in{\mathcal{F}}_{K}$. To avoid distracting technicalities, we assume that the material properties are only mildly heterogeneous and anisotropic. Thus, we hide anisotropy ratios such as $\alpha^{\sharp}_{K}/\alpha^{\flat}_{K}$ in the generic constants used in the error analysis. We do the same for contrast factors such as $\alpha^{\sharp}_{K}/\alpha^{\sharp}_{K^{\prime}}$, where $K^{\prime}$ lies in some neighborhood of $K$. We refer the reader to Remarks 6.25 and 6.28 for some further insight.
@@ 6.4 $hp$-a posteriori error analysis
Invoke the triangle inequality together with Lemmas 6.30 and 6.31, and recall the definition of $\eta$ from (6.13). ∎
@@
Local bounds on all terms composing the error estimator $\eta_{K}$ can be found in [12, Theorem 6.5] except for the term involving ${\nabla_{h}{\times}{\boldsymbol{E}}_{h}}]\!]_{F}^{\mathrm{d}}$. Invoking the $H^{1}$- to $L^{2}$-norm inverse inequality from [39, Theorem 4.76], we infer that